OpenStudy (aravindg):

A 100 k/h wind blows normally against one wall of a house with an area of 50 m square.calculate the force exerted on the wall if the air moves parallel to wall after striking it and has a density of 1.134 kg /m cube

OpenStudy (aravindg):

help plzzzzzzzzzzz

OpenStudy (anonymous):

not worrying about units q= pressure p= density v= velocity q= 1/2 p v^2 q = 1/2 1.134 100^2 5670 is pressure I guess to get the force we multiply by area 5670*50=

OpenStudy (aravindg):

r u sure???

OpenStudy (anonymous):

about 80 percent

OpenStudy (aravindg):

rest too plz

OpenStudy (anonymous):

No, I meant I am 80% sure